18. 3/4 points | Previous Answers PODStat5 7.E.069 My Notes Ask Your Determine the value of z such that it satisfies the conditions below. (Round your answers to two decimal places.) (a)-z* and z* separate the middle 94% of all z values from the most extreme 696. z = 11.88 (b)-z* and z* separate the middle 90% of all z values from the most extreme 10% z 1.64 (c)-z* and z* separate the middle 98% of all z values from the most extreme 296. z-2.30 (d)-z and z* separate the middle 92% of all z values from the most extreme 896. z* 175 You may need to use the appropriate table in Appendix A to answer this question. Need Help? Read ItTalk to a TutorDetermine the value of z* such that it satisfies the conditions below. (Round your answers to two decimal places.)
Answer #1

18) Solution :

c) Solution :
P(-z < Z < z) = 0.98
P(Z < z) - P(Z < z) = 0.98
2P(Z < z) - 1 = 0.98
2P(Z < z) = 1 + 0.98
2P(Z < z) = 1.98
P(Z < z) = 1.98 / 2 = 0.99
P(Z < 2.33) = 0.99
z* = 2.33